\providecommand{\href}[2]{#2}\begingroup\raggedright\begin{thebibliography}{10}

\bibitem{Liu:1987nz}
J.~Liu and J.~Polchinski, \emph{{Renormalization of the Mobius Volume}},
  \href{https://doi.org/10.1016/0370-2693(88)91566-3}{\emph{Phys. Lett. B}
  {\bfseries 203} (1988) 39}.

\bibitem{Maldacena:2001km}
J.~M. Maldacena and H.~Ooguri, \emph{{Strings in ${\rm AdS}_3$ and ${\rm
  SL}(2,\mathds{R})$ WZW model. Part 3. Correlation functions}},
  \href{https://doi.org/10.1103/PhysRevD.65.106006}{\emph{Phys. Rev.}
  {\bfseries D65} (2002) 106006}
  [\href{https://arxiv.org/abs/hep-th/0111180}{{\ttfamily hep-th/0111180}}].

\bibitem{Erbin:2019uiz}
H.~Erbin, J.~Maldacena and D.~Skliros, \emph{{Two-Point String Amplitudes}},
  \href{https://doi.org/10.1007/JHEP07(2019)139}{\emph{JHEP} {\bfseries 07}
  (2019) 139} [\href{https://arxiv.org/abs/1906.06051}{{\ttfamily
  1906.06051}}].

\bibitem{Eberhardt:2021ynh}
L.~Eberhardt and S.~Pal, \emph{{The disk partition function in string theory}},
  \href{https://doi.org/10.1007/JHEP08(2021)026}{\emph{JHEP} {\bfseries 08}
  (2021) 026} [\href{https://arxiv.org/abs/2105.08726}{{\ttfamily
  2105.08726}}].

\bibitem{Gibbons:1976ue}
G.~W. Gibbons and S.~W. Hawking, \emph{{Action Integrals and Partition
  Functions in Quantum Gravity}},
  \href{https://doi.org/10.1103/PhysRevD.15.2752}{\emph{Phys. Rev. D}
  {\bfseries 15} (1977) 2752}.

\bibitem{Mahajan:2021nsd}
R.~Mahajan, D.~Stanford and C.~Yan, \emph{{Sphere and disk partition functions
  in Liouville and in matrix integrals}},
  \href{https://doi.org/10.1007/JHEP07(2022)132}{\emph{JHEP} {\bfseries 07}
  (2022) 132} [\href{https://arxiv.org/abs/2107.01172}{{\ttfamily
  2107.01172}}].

\bibitem{Anninos:2021ene}
D.~Anninos, T.~Bautista and B.~M\"uhlmann, \emph{{The two-sphere partition
  function in two-dimensional quantum gravity}},
  \href{https://doi.org/10.1007/JHEP09(2021)116}{\emph{JHEP} {\bfseries 09}
  (2021) 116} [\href{https://arxiv.org/abs/2106.01665}{{\ttfamily
  2106.01665}}].

\bibitem{Erler:2022agw}
T.~Erler, \emph{{The closed string field theory action vanishes}},
  \href{https://doi.org/10.1007/JHEP10(2022)055}{\emph{JHEP} {\bfseries 10}
  (2022) 055} [\href{https://arxiv.org/abs/2204.12863}{{\ttfamily
  2204.12863}}].

\bibitem{Tseytlin:1988tv}
A.~A. Tseytlin, \emph{{Mobius Infinity Subtraction and Effective Action in
  $\sigma$ Model Approach to Closed String Theory}},
  \href{https://doi.org/10.1016/0370-2693(88)90421-2}{\emph{Phys. Lett. B}
  {\bfseries 208} (1988) 221}.

\bibitem{Ahmadain:2022tew}
A.~Ahmadain and A.~C. Wall, \emph{{Off-Shell Strings I: S-matrix and Action}},
  \href{https://arxiv.org/abs/2211.08607}{{\ttfamily 2211.08607}}.

\bibitem{Maldacena:1997re}
J.~M. Maldacena, \emph{{The Large $N$ limit of superconformal field theories
  and supergravity}}, \href{https://doi.org/10.1023/A:1026654312961,
  10.4310/ATMP.1998.v2.n2.a1}{\emph{Int. J. Theor. Phys.} {\bfseries 38} (1999)
  1113} [\href{https://arxiv.org/abs/hep-th/9711200}{{\ttfamily
  hep-th/9711200}}].

\bibitem{Eberhardt:2018ouy}
L.~Eberhardt, M.~R. Gaberdiel and R.~Gopakumar, \emph{{The Worldsheet Dual of
  the Symmetric Product CFT}},
  \href{https://doi.org/10.1007/JHEP04(2019)103}{\emph{JHEP} {\bfseries 04}
  (2019) 103} [\href{https://arxiv.org/abs/1812.01007}{{\ttfamily
  1812.01007}}].

\bibitem{Eberhardt:2021vsx}
L.~Eberhardt, \emph{{A perturbative CFT dual for pure NS-NS $\mathrm{AdS}_{3}$
  strings}}, \href{https://doi.org/10.1088/1751-8121/ac47b2}{\emph{J. Phys. A}
  {\bfseries 55} (2022) 064001}
  [\href{https://arxiv.org/abs/2110.07535}{{\ttfamily 2110.07535}}].

\bibitem{Henningson:1998gx}
M.~Henningson and K.~Skenderis, \emph{{The Holographic Weyl anomaly}},
  \href{https://doi.org/10.1088/1126-6708/1998/07/023}{\emph{JHEP} {\bfseries
  07} (1998) 023} [\href{https://arxiv.org/abs/hep-th/9806087}{{\ttfamily
  hep-th/9806087}}].

\bibitem{Troost:2011ud}
J.~Troost, \emph{{The $\mathrm{AdS}_3$ central charge in string theory}},
  \href{https://doi.org/10.1016/j.physletb.2011.10.007}{\emph{Phys. Lett. B}
  {\bfseries 705} (2011) 260}
  [\href{https://arxiv.org/abs/1109.1923}{{\ttfamily 1109.1923}}].

\bibitem{Giveon:1998ns}
A.~Giveon, D.~Kutasov and N.~Seiberg, \emph{{Comments on string theory on ${\rm
  AdS}_3$}}, \href{https://doi.org/10.4310/ATMP.1998.v2.n4.a3}{\emph{Adv.
  Theor. Math. Phys.} {\bfseries 2} (1998) 733}
  [\href{https://arxiv.org/abs/hep-th/9806194}{{\ttfamily hep-th/9806194}}].

\bibitem{Polchinski:1998rq}
J.~Polchinski, \emph{{String theory. Vol. 1: An introduction to the bosonic
  string}}, Cambridge Monographs on Mathematical Physics. Cambridge University
  Press, 12, 2007,
  \href{https://doi.org/10.1017/CBO9780511816079}{10.1017/CBO9780511816079}.

\bibitem{Knizhnik:1988ak}
V.~G. Knizhnik, A.~M. Polyakov and A.~B. Zamolodchikov, \emph{{Fractal
  Structure of 2D Quantum Gravity}},
  \href{https://doi.org/10.1142/S0217732388000982}{\emph{Mod. Phys. Lett. A}
  {\bfseries 3} (1988) 819}.

\bibitem{Gawedzki:1991yu}
K.~Gawedzki, \emph{{Noncompact WZW Conformal Field Theories}},  in \emph{{Nato
  Advanced Study Institute: New Symmetry Principles in Quantum Field Theory}},
  pp.~0247--274, 10, 1991,
  \href{https://arxiv.org/abs/hep-th/9110076}{{\ttfamily hep-th/9110076}}.

\bibitem{Ishibashi:2000fn}
N.~Ishibashi, K.~Okuyama and Y.~Satoh, \emph{{Path integral approach to string
  theory on $\mathrm{AdS}_3$}},
  \href{https://doi.org/10.1016/S0550-3213(00)00509-5}{\emph{Nucl. Phys. B}
  {\bfseries 588} (2000) 149}
  [\href{https://arxiv.org/abs/hep-th/0005152}{{\ttfamily hep-th/0005152}}].

\bibitem{Hosomichi:2000bm}
K.~Hosomichi, K.~Okuyama and Y.~Satoh, \emph{{Free field approach to string
  theory on $\mathrm{AdS}_3$}},
  \href{https://doi.org/10.1016/S0550-3213(01)00002-5}{\emph{Nucl. Phys. B}
  {\bfseries 598} (2001) 451}
  [\href{https://arxiv.org/abs/hep-th/0009107}{{\ttfamily hep-th/0009107}}].

\bibitem{Eberhardt:2019ywk}
L.~Eberhardt, M.~R. Gaberdiel and R.~Gopakumar, \emph{{Deriving the
  $\mathrm{AdS}_{3}/\mathrm{CFT}_{2}$ correspondence}},
  \href{https://doi.org/10.1007/JHEP02(2020)136}{\emph{JHEP} {\bfseries 02}
  (2020) 136} [\href{https://arxiv.org/abs/1911.00378}{{\ttfamily
  1911.00378}}].

\bibitem{Dei:2021yom}
A.~Dei and L.~Eberhardt, \emph{{String correlators on $\mathrm{AdS}_{3}$:
  four-point functions}},
  \href{https://doi.org/10.1007/JHEP09(2021)209}{\emph{JHEP} {\bfseries 09}
  (2021) 209} [\href{https://arxiv.org/abs/2107.01481}{{\ttfamily
  2107.01481}}].

\bibitem{Dei:2022pkr}
A.~Dei and L.~Eberhardt, \emph{{String correlators on $\mathrm{AdS}_3$:
  Analytic structure and dual CFT}},
  \href{https://doi.org/10.21468/SciPostPhys.13.3.053}{\emph{SciPost Phys.}
  {\bfseries 13} (2022) 053}
  [\href{https://arxiv.org/abs/2203.13264}{{\ttfamily 2203.13264}}].

\bibitem{Kutasov:1999xu}
D.~Kutasov and N.~Seiberg, \emph{{More comments on string theory on ${\rm
  AdS}_3$}}, \href{https://doi.org/10.1088/1126-6708/1999/04/008}{\emph{JHEP}
  {\bfseries 04} (1999) 008}
  [\href{https://arxiv.org/abs/hep-th/9903219}{{\ttfamily hep-th/9903219}}].

\bibitem{Kim:2015gak}
J.~Kim and M.~Porrati, \emph{{On the central charge of spacetime current
  algebras and correlators in string theory on $\mathrm{AdS}_{3}$}},
  \href{https://doi.org/10.1007/JHEP05(2015)076}{\emph{JHEP} {\bfseries 05}
  (2015) 076} [\href{https://arxiv.org/abs/1503.07186}{{\ttfamily
  1503.07186}}].

\bibitem{Giveon:2001up}
A.~Giveon and D.~Kutasov, \emph{{Notes on $\mathrm{AdS}_3$}},
  \href{https://doi.org/10.1016/S0550-3213(01)00573-9}{\emph{Nucl. Phys. B}
  {\bfseries 621} (2002) 303}
  [\href{https://arxiv.org/abs/hep-th/0106004}{{\ttfamily hep-th/0106004}}].

\bibitem{Gerchkovitz:2014gta}
E.~Gerchkovitz, J.~Gomis and Z.~Komargodski, \emph{{Sphere Partition Functions
  and the Zamolodchikov Metric}},
  \href{https://doi.org/10.1007/JHEP11(2014)001}{\emph{JHEP} {\bfseries 11}
  (2014) 001} [\href{https://arxiv.org/abs/1405.7271}{{\ttfamily 1405.7271}}].

\bibitem{Skenderis:2002wp}
K.~Skenderis, \emph{{Lecture notes on holographic renormalization}},
  \href{https://doi.org/10.1088/0264-9381/19/22/306}{\emph{Class. Quant. Grav.}
  {\bfseries 19} (2002) 5849}
  [\href{https://arxiv.org/abs/hep-th/0209067}{{\ttfamily hep-th/0209067}}].

\bibitem{Brown:1986nw}
J.~Brown and M.~Henneaux, \emph{{Central Charges in the Canonical Realization
  of Asymptotic Symmetries: an Example from Three-Dimensional Gravity}},
  \href{https://doi.org/10.1007/BF01211590}{\emph{Commun. Math. Phys.}
  {\bfseries 104} (1986) 207}.

\bibitem{Beccaria:2014qea}
M.~Beccaria, G.~Macorini and A.~A. Tseytlin, \emph{{Supergravity one-loop
  corrections on $\mathrm{AdS}_7$ and $\mathrm{AdS}_3$, higher spins and
  AdS/CFT}}, \href{https://doi.org/10.1016/j.nuclphysb.2015.01.014}{\emph{Nucl.
  Phys. B} {\bfseries 892} (2015) 211}
  [\href{https://arxiv.org/abs/1412.0489}{{\ttfamily 1412.0489}}].

\bibitem{Callan:1991at}
C.~G. Callan, Jr., J.~A. Harvey and A.~Strominger, \emph{{Supersymmetric string
  solitons}},  \href{https://arxiv.org/abs/hep-th/9112030}{{\ttfamily
  hep-th/9112030}}.

\bibitem{Giveon:1999px}
A.~Giveon and D.~Kutasov, \emph{{Little string theory in a double scaling
  limit}}, \href{https://doi.org/10.1088/1126-6708/1999/10/034}{\emph{JHEP}
  {\bfseries 10} (1999) 034}
  [\href{https://arxiv.org/abs/hep-th/9909110}{{\ttfamily hep-th/9909110}}].

\bibitem{Witten:2012bh}
E.~Witten, \emph{{Superstring Perturbation Theory Revisited}},
  \href{https://arxiv.org/abs/1209.5461}{{\ttfamily 1209.5461}}.

\bibitem{Blumenhagen:2013fgp}
R.~Blumenhagen, D.~L\"ust and S.~Theisen, \emph{{Basic concepts of string
  theory}}, Theoretical and Mathematical Physics. Springer, Heidelberg,
  Germany, 2013,
  \href{https://doi.org/10.1007/978-3-642-29497-6}{10.1007/978-3-642-29497-6}.

\bibitem{Teschner:1997ft}
J.~Teschner, \emph{{On structure constants and fusion rules in the
  $\mathrm{SL}(2,\mathbb{C}) / \mathrm{SU}(2)$ WZNW model}},
  \href{https://doi.org/10.1016/S0550-3213(99)00072-3}{\emph{Nucl. Phys. B}
  {\bfseries 546} (1999) 390}
  [\href{https://arxiv.org/abs/hep-th/9712256}{{\ttfamily hep-th/9712256}}].

\bibitem{Dei:2021xgh}
A.~Dei and L.~Eberhardt, \emph{{String correlators on $\mathrm{AdS}_{3}$:
  three-point functions}},
  \href{https://doi.org/10.1007/JHEP08(2021)025}{\emph{JHEP} {\bfseries 08}
  (2021) 025} [\href{https://arxiv.org/abs/2105.12130}{{\ttfamily
  2105.12130}}].

\bibitem{Maldacena:2000hw}
J.~M. Maldacena and H.~Ooguri, \emph{{Strings in ${\rm AdS}_3$ and ${\rm
  SL}(2,\mathds{R})$ WZW model 1.: The Spectrum}},
  \href{https://doi.org/10.1063/1.1377273}{\emph{J. Math. Phys.} {\bfseries 42}
  (2001) 2929} [\href{https://arxiv.org/abs/hep-th/0001053}{{\ttfamily
  hep-th/0001053}}].

\end{thebibliography}\endgroup

\documentclass[12pt,a4paper]{article}
\usepackage{jheppub}  
\usepackage{amssymb} 
\usepackage{amsmath}
\usepackage{mathtools}
\usepackage{amsfonts}    
\usepackage{dsfont}
\usepackage{pdfpages}
\usepackage{verbatim}
\usepackage{tensor}
\usepackage{mathrsfs}
\usepackage[mathscr]{euscript}
\usepackage{tikz}
\usetikzlibrary{patterns,decorations.pathreplacing}
\usepackage{slashed}

\newcommand{\ba}{\begin{align}}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\def\bd{\begin{tikzpicture}}
\def\ed{\end{tikzpicture}}

\usepackage{tcolorbox}

\renewcommand\d{\text{d}}

\newcommand\ZZ{\mathbb{Z}}
\newcommand\RR{\mathbb{R}}

\newcommand\SL{\text{SL}}

\allowdisplaybreaks[1]

\title{Holographic Weyl anomaly in string theory}

\author[1]{Lorenz Eberhardt} 
\author[1,2]{\!\!, Sridip Pal} 
\affiliation[1]{School of Natural Sciences, Institute for Advanced Study, 1 Einstein Drive,
Princeton,  NJ 08540, USA}
\affiliation[2]{Walter Burke Institute for Theoretical Physics,  California Institute of Technology,  Pasadena, CA, USA}
\emailAdd{elorenz@ias.edu}
\emailAdd{sridip@caltech.edu}

\abstract{
We compute the worldsheet sphere partition function of string theory on global AdS$_3$ with pure NS-NS flux.
Because of an unfixed M\"obius symmetry on the worldsheet, there is a cancellation of infinities and only a part of the answer is unambiguous. We show that it precisely reproduces the holographic Weyl anomaly and the ambiguous terms correspond to the possible counterterms of the boundary CFT.
}

\begin{document}

\maketitle

\makeatletter
\g@addto@macro\bfseries{\boldmath}
\makeatother

\section{Introduction}
The seemingly simplest string theory diagrams are paradoxically under poorest conceptual control since they often reflect ambiguities of the theory.
This happens when the worldsheet has a residual non-compact automorphism group. Since it is part of the gauge group in string theory, the infinite volume needs to be canceled by a corresponding target space volume divergence. The possible topologies with a non-compact automorphism group are the sphere with $0$, $1$ or $2$ punctures and the disk with $0$, $1$ or $2$ boundary punctures. The disk case and the two-point function on the sphere can be treated without ambiguities for example by imposing a suitable gauge-fixing condition \cite{Liu:1987nz,Maldacena:2001km, Erbin:2019uiz, Eberhardt:2021ynh}. 
However, this leaves out the arguably most interesting case -- the sphere partition function.

The sphere partition function conjecturally captures the on-shell gravitational action of the background and as such carries the usual ambiguities \cite{Gibbons:1976ue}. Using standard worldsheet techniques, it has only been computed directly in some very special situations such as the minimal string \cite{Mahajan:2021nsd} or in two-dimensional gravity \cite{Anninos:2021ene}. For compact target spaces, the divergence from the infinite volume of the automorphism group is uncompensated and the sphere partition function vanishes \cite{Erler:2022agw}.
It has also been proposed that it might be necessary to give up conformality on the worldsheet to compute the sphere partition function in general \cite{Tseytlin:1988tv, Ahmadain:2022tew}.

We explain the physically relevant example of $\mathrm{AdS}_3$ backgrounds where the sphere partition function can be directly understood within the standard framework of string perturbation theory. Via the AdS/CFT correspondence, these string backgrounds are dual to two-dimensional CFTs \cite{Maldacena:1997re, Eberhardt:2018ouy, Eberhardt:2021vsx}. The sphere partition function captures the conformal anomaly in the boundary CFT and reflects the process of holographic renormalization in string theory \cite{Henningson:1998gx}. Contrary to ordinary gravity, we don't have to do any adhoc regularization of the gravitational on-shell action.\footnote{\cite{Troost:2011ud} computed the one-point function on $\mathrm{AdS}_3$ using a certain cutoff procedure. Hence it might also be possible to implement such a procedure for the sphere partition function itself.}

The method of our computation is not new -- we exploit that we can compute derivatives of the sphere partition function by inserting zero-momentum dilaton vertex operators which we explain in Section~\ref{sec:computing}. The interpretation of the result is however interesting and we discuss it in detail in Section~\ref{sec:interpreting}. For the actual computation, we need to carefully work out the normalization of the sphere partition function, which we do in Appendix~\ref{app:normalization}. 

\section{Computing the sphere partition function} \label{sec:computing}
We consider superstrings on $\text{AdS}_3 \times X$ with pure NS-NS flux. We are agnostic about the compactification $X$ since it will not play a role. We are only interested in genus $0$ and thus do not have to specify the GSO projection. We merely have to assume that $X$ is a compact sigma model described by an $\mathcal{N}=1$ sigma-model on the worldsheet of the correct central charge. The $\mathrm{AdS}_3$ part of the background is described by the $\SL(2,\RR)$ WZW model at level $k+2$, together with three free fermions \cite{Giveon:1998ns}. The most important part of the worldsheet theory is the $\SL(2,\RR)$ WZW model whose relevant features we briefly recall (or more precisely we are actually considering an analytic continuation of the WZW model to Euclidean spacetime signature known as the $H_3^+$ model). The fermions do not influence our computations, except for shifting the level $k \to k+2$ and setting the correct normalization for the dilaton vertex operator.
\subsection{The action}
The action of the $\SL(2,\RR)_{k+2}$ model reads\footnote{Since numerical factors will be important, we notice that we define $\d^2z=\d x\, \d y$ and $\partial=\frac{1}{2}(\partial_x-i \partial y)$ for $z=x+i y$. We follow the conventions for sigma-models given e.g.\ in Polchinski \cite[Section 3.7]{Polchinski:1998rq} with $\alpha'=1$.}
\be 
S_{\SL(2,\RR)}=\frac{k+2}{\pi}\int \d^2 z \, \big(\partial \phi \bar{\partial} \phi+\mathrm{e}^{2\phi} \partial \bar{\gamma} \bar{\partial}\gamma\big)\ .
\ee
This is simply the action for a sigma model on Euclidean global $\text{AdS}_3$   in Poincar\'e coordinates.\footnote{We take $\bar{\gamma}$ to be the complex conjugate of $\gamma$. For Lorentzian $\mathrm{AdS}_3$, they would both be real and independent. Thus we are considering an analytic continuation of the $\SL(2,\RR)$ WZW model. As is common in the literature, we will nonetheless continue to refer to this CFT as the $\SL(2,\RR)$ WZW model.} The metric in these coordinates reads
\be 
\d s^2_\text{AdS}=\d \phi^2+\mathrm{e}^{2\phi} \, \d \gamma \, \d \bar{\gamma}\ . \label{eq:AdS3 metric}
\ee
The boundary of AdS$_3$ is located at $\phi \to \infty$.\footnote{The change of coordinates $z=\mathrm{e}^{-\phi}$ gives a perhaps more standard form of the Poincar\'e metric.} We use units in which $\alpha'=1$ so that $k$ corresponds to the radius of AdS$_3$ measured in units of string length.
The $B$-field $B=k\, \d\gamma \wedge d \bar{\gamma}$ is responsible for the absence of the term $\mathrm{e}^{2\phi}\partial \gamma \bar{\partial} \bar{\gamma}$ in the action. The shift $k \to k+2$ is a one-loop effect and thus we omitted it in semiclassical quantities.

It is convenient to modify the action slightly and introduce a coupling constant $\mu$ analogous to the cosmological constant in Liouville theory:
\be 
S_{\SL(2,\RR)}=\frac{1}{4\pi}\int \d^2 z \ \big(4(k+2)\partial \phi \bar{\partial} \phi+\mu\mathrm{e}^{2\phi} \partial \bar{\gamma} \bar{\partial}\gamma\big) \label{eq:SL2R action}
\ee
This seems to be unimportant since we can always remove $\mu$ by redefining $\phi \to \phi-\frac{1}{2} \log \mu$. However, the correlation functions and partition functions will not be quite invariant under this shift since both the vertex operators and the path integral measure are not invariant. This leads to a version of the KPZ scaling argument \cite{Knizhnik:1988ak}. The path integral measure takes the form
\be 
\mathscr{D}\phi \, \mathscr{D}(\mathrm{e}^\phi \gamma)\, \mathscr{D}(\mathrm{e}^\phi \bar{\gamma})\ ,
\ee
which is induced from the Haar measure on $\SL(2,\RR)$. 
One can remove the various factors of $\mathrm{e}^\phi$ at the cost of introducing a Jacobian factor which corrects the action at one loop. The result is \cite{Gawedzki:1991yu, Giveon:1998ns, Ishibashi:2000fn, Hosomichi:2000bm} 
\be 
S=\frac{1}{4\pi} \int \d^2 z \, \big( 4k \partial \phi \bar{\partial} \phi+R \phi+\mu  \mathrm{e}^{2\phi} \partial \bar{\gamma} \bar{\partial} \gamma \big)\ , \label{eq:SL2R action quantum phi gamma}
\ee
where the path integral measure is now that of a free fields. This result can be derived in a variety of ways, for example by (i) by relating it to the chiral anomaly as in \cite{Gawedzki:1991yu} and using the well-known formulas for the chiral anomaly in 2d, (ii) by using an index theorem counting the zero modes of $\gamma$ (and $\beta$ that is introduced to pass to a first order formalism), or (iii) by simply writing down the most general local expression of the correct dimension in $\phi$ and the metric, requiring the Jacobian to behave group-like under repeated removal of factors $\mathrm{e}^\phi$ and matching the central charge to the known value of $\mathrm{SL}(2,\mathbb{R})_{k+2}$.

One can see from \eqref{eq:SL2R action quantum phi gamma} that the path integral over $\mathrm{e}^{-S}$ converges for surfaces of negative Euler characteristics. Indeed, in this case the path integral is damped for $\phi \to +\infty$ thanks to the presence of the term $\mu \mathrm{e}^{2\phi} \partial \bar{\gamma} \bar{\partial} \gamma$ and for $\phi \to -\infty$ thanks to the presence of the Ricci scalar.\footnote{The suppression for $\phi \to -\infty$ can fail on a genus 0 surface or when including certain vertex operators. This leads to singularities in the correlators as a function of the spins.
The suppression for $\phi \to +\infty$ can also fail because there might be a holomorphic map $\gamma$ with $\bar{\partial}\gamma=0$. This leads to certain well-studied singularities in the correlation functions of the model as a function of the moduli \cite{Maldacena:2001km, Eberhardt:2019ywk, Dei:2021yom, Dei:2022pkr}.}

\subsection{KPZ scaling and the coupling constant}
From the form \eqref{eq:SL2R action quantum phi gamma}, it is simple to derive the $\mu$-dependence of a partition function. We can shift $\phi \to \phi-\frac{1}{2} \log \mu$. Only the term involving the Ricci scalar is not fully invariant under this shift and shows that the partition function has the $\mu$-dependence $\mu^{\frac{1}{2}\chi_g}=\mu^{1-g}$ as a consequence of the Gauss-Bonnet theorem. More generally, in the presence of $n$ vertex operators of $\SL(2,\RR)$ spin $j_i$, we get the following scaling of a correlation function:
\be 
\bigg \langle \prod_i V_{j_i} \bigg \rangle \propto \mu^{1-g-\sum_i j_i}\ . \label{eq:KPZ scaling}
\ee
This is the analogue of the KPZ scaling of Liouville theory \cite{Knizhnik:1988ak}. 
We should note that this equation only follows from the path integral if the exponent of $\mu$ is negative, since otherwise the path integral does not converge. We will indeed see below that \eqref{eq:KPZ scaling} can get modified for positive exponents of $\mu$.

We also note that $\mu$ plays the role of the string coupling in the theory, more precisely, we can identify
\be 
\mu\propto \frac{1}{g_\text{s}^2}\ .
\ee
In particular, there is no need to introduce a separate string coupling since it is already contained naturally in the $\SL(2,\RR)_{k+2}$ WZW model. In particular, weak string coupling corresponds to large values of $\mu$.
\subsection{The sphere partition function and its derivatives}
We are interested in computing the sphere partition function of superstrings on $\text{AdS}_3 \times X$. As usual, this computation is subtle since one has to divide by the volume of the super M\"obius group $\text{OSP}(1|2,\mathbb{C})/\ZZ_2$, which is badly divergent \cite{Liu:1987nz}.
Instead, one can use that insertion of the vertex operator
\be 
I=-\frac{1}{4\pi}\, \mathrm{e}^{2\phi} \partial \bar{\gamma}\bar{\partial}\gamma \label{eq:definition I}
\ee
implements $\mu$-derivatives of correlation functions. One can read off from the exponent that it indeed has spin $j=1$ consistent with the scaling \eqref{eq:KPZ scaling}. 
This is the zero mode of the dilaton vertex operator and has played a prominent role in $\mathrm{AdS}_3$ holography \cite{Giveon:1998ns, Kutasov:1999xu, Kim:2015gak, Eberhardt:2021vsx}.
From a path integral point of view we have for the $\SL(2,\RR)_{k+2}$ WZW model 
\be 
\partial_\mu \, \bigg \langle \prod_i V_{j_i}(x_i, z_i) \bigg \rangle_{\!\! g}=\int \mathrm{d}^2 z \ \bigg \langle I(z) \prod_i V_{j_i}(x_i,z_i) \bigg \rangle_{\!\! g}\ , \label{eq:mu derivative correlator}
\ee
since the dilaton vertex operator is precisely the marginal operator appearing with the coupling $\mu$ in eq.~\eqref{eq:SL2R action quantum phi gamma}.
Here we use the natural primary vertex operators $V_{j_i}(x_i,z_i)$ of the sigma-model on Euclidean $\text{AdS}_3$. We note that the coordinate $x$ corresponds to the location of the vertex operator on the boundary of Euclidean AdS$_3$.
In view of the scaling \eqref{eq:KPZ scaling}, the left-hand side of \eqref{eq:mu derivative correlator} is simply the correlation function itself times the exponent $1-g-\sum_i j_i$ and times $\mu^{-1}$. This equation can also be established directly from an axiomatic approach to the $\SL(2,\RR)$ WZW model \cite{Giveon:2001up, Kim:2015gak}. 

On the level of the integrated string theory correlators, this means that
\be 
\partial_\mu \, \bigg\langle\!\!\!\bigg\langle  \prod_i V_{j_i}(x_i)\bigg \rangle\!\!\!\bigg \rangle_{\!\! g}=\bigg\langle\!\!\!\bigg\langle  I\prod_i V_{j_i}(x_i)\bigg \rangle\!\!\!\bigg \rangle_{\!\! g}\ , \label{eq:mu derivative I insertion string theory}
\ee
where the double brackets denote the integrated string theory correlators, which no longer depend on $z_i$.\footnote{For arbitrary choices of spins $j_i$, one may need to combine the vertex operators with vertex operator from the internal CFT $X$ so that they satisfy the mass-shell condition. This won't be important for us.} Since we are working within the RNS formalism of superstring theory, we should also include picture changing operators or integrate over supermoduli space. We use the former formalism, but suppress this from our notation in the main text, where we just explain the bosonic formalism. We include the details in the computation of Appendix~\ref{app:normalization}.

The starting point for the sphere partition function is not a well-defined one, but it is reasonable to assume that it can be defined by assuming \eqref{eq:mu derivative I insertion string theory} to continue to hold.
In particular, we learn that for the string theory sphere partition function $Z_{\mathrm{S}^2}$  (which could also be represented by an empty double bracket), we can compute the third derivative without problems:
\be 
\partial_\mu^3 Z_{\mathrm{S}^2}=\langle\!\langle I \, I\, I \rangle \! \rangle=C_{\text{S}^2} \langle I(0) I(1) I(\infty) \rangle\ ,
\ee
since a string theory three-point function essentially coincides with the worldsheet three-point function up to ghosts and the normalization $C_{\text{S}^2}$ of the string theory path integral. The right hand side is a well-defined quantity in the $\SL(2,\RR)$ WZW model and can be computed. It takes the form
\be 
\partial_\mu^3 Z_{\mathrm{S}^2}=\langle\!\langle I\, I \, I\rangle\!\rangle=-F(k)\, \mu^{-2} \label{eq:third derivative sphere}
\ee
for some function $F(k)$, which we compute in Appendix~\ref{app:normalization}. We used the KPZ scaling \eqref{eq:KPZ scaling} for the $\mu$ dependence. We can thus integrate this equation back up to obtain the sphere partition function, up to three integration constants. 

We could have done slightly better. It is known how to define the two-point function in string theory, see \cite{Maldacena:2001km} for the case of $\text{AdS}_3$ and \cite{Erbin:2019uiz} for a more general discussion. Thus one can reliably compute the second derivative of $Z_{\mathrm{S}^2}$ and obtains
\be 
\partial_\mu^2 Z_{\mathrm{S}^2}=\langle\!\langle I\, I \rangle\!\rangle=F(k) \, \mu^{-1}\ . \label{eq:second derivative sphere}
\ee
This means that there is no integration constant when integrating \eqref{eq:third derivative sphere} back up to \eqref{eq:second derivative sphere}. We thus obtain
\be 
Z_{\mathrm{S}^2}=F(k) \, \mu \log \mu + C_1 +C_2 \mu \label{eq:sphere partition function integration constants}
\ee
for two integration constants $C_1$ and $C_2$. We notice that they should be included since there is no natural scale in the logarithm and thus the first well-defined term can mix with the integration constants.

\section{Interpreting the result} \label{sec:interpreting}
We now interpret the result \eqref{eq:sphere partition function integration constants} physically.
\subsection{Expectation}
Let us first review what result is expected. From a holography point of view, our computation should reproduce the large $N$ partition function of the dual CFT on the boundary of global Euclidean $\mathrm{AdS}_3$, which also happens to be a two-sphere (not to be confused with the worldsheet two-sphere that we discussed above). The sphere partition function of a CFT is fully determined from the conformal anomaly. Assuming that the metric is a round sphere of radius $R$, it takes the form
\be 
\log Z_\text{CFT}=\frac{c}{3} \log R+N_1+N_2 R^2\ , \label{eq:dual CFT sphere partition function}
\ee
where $c$ is the central charge. The constants $N_1$ and $N_2$ correspond to the two possible counterterms we can add to the theory. $N_1$ and $N_2$ arise since we can add the counterterms
\be 
\int \mathrm{d}^2 x \, \sqrt{g} R\quad \text{and}\quad 
\int \mathrm{d}^2x\, \sqrt{g}
\ee
to the action, which modifies the partition function accordingly. For a nice general discussion about these ambiguities, see e.g.\ \cite{Gerchkovitz:2014gta}. Of course, it is well-known how to reproduce this structure from classical gravity via holographic renormalization \cite{Henningson:1998gx, Skenderis:2002wp}.

The sphere partition function in string theory should hence precisely compute this universal dependence of the partition function on the radius and reflect the possible ambiguities of the sphere partition function of the dual CFT. More concretely, we should have
\be 
Z_{\mathrm{S}^2}+\mathcal{O}(\log \mu)\overset{!}{=} \log Z_\text{CFT}\ .
\ee
The logarithm on the right hand side appears since the full string partition function also receives contributions from several disconnected spheres. The $\log \mu$ corrections on the string side originate from the torus diagram, see Section~\ref{subsec:torus partition function} below.

\subsection{Relating the coupling \texorpdfstring{$\mu$}{mu} to the radius}
Going back to the string theory calculation, we notice that the expected result \eqref{eq:dual CFT sphere partition function} features an explicit dependence on the radius of the boundary sphere on which the dual CFT is defined, which of course is not a parameter that entered the string theory calculation.

From the path integral definition \eqref{eq:SL2R action} of the $\SL(2,\RR)$ action one can however plausibly relate the coupling $\mu$ to the size of the asymptotic boundary as follows. In the gravity calculation, we would put the AdS boundary explicitly on some profile $\phi =\phi(\gamma,\bar{\gamma})$ and $\phi$ would become the Weyl factor of the induced metric. 
With the inclusion of the coupling constant $\mu$, the induced metric is instead
\be 
\mu\,  \mathrm{e}^{2\phi} \, \d \gamma\, \d \bar{\gamma}\ .
\ee
In particular, we see that $\mu$ is related to the size of the metric on the holographic surface. We can thus identify
\be 
\mu \propto R^2\ , \label{eq:mu R relation}
\ee
where $R$ is the characteristic radius of the holographic surface. 

Hence we see that we should interpret some of the $\mu$'s in \eqref{eq:sphere partition function integration constants} in terms of the radius of the holographic surface. We write
\be 
Z_{\mathrm{S}^2}=2 F(k) \mu \log R+ C_1+C_2' \mu+ C_2'' R^2\ . \label{eq:string sphere partition function mu R replaced}
\ee
Intuitively, we should trade the non-analytic dependence on $\mu$'s that does not follow the expected naive scaling \eqref{eq:KPZ scaling} for $R$'s. For the term proportional to $\mu$, it does not matter whether we interpret $\mu$ as $R^2$ and we included both terms. In any case, this has the correct form to match with \eqref{eq:dual CFT sphere partition function}, where we keep in mind that e.g.\ the central charge scales like $\mu \propto \frac{1}{g_\text{s}^2}\propto G_\text{N}^{-1}$ \cite{Brown:1986nw}.

Besides having the correct form, the only meaningful comparison arises from comparing the precise coefficient of $\log R$.
To claim success, we hence have to check that
\be 
6F(k) \mu= c \label{eq:F(k) c relation}
\ee
reproduces the expected central charge $c$ of the boundary CFT. 

Let us outline the strategy for this. In order to relate $F(k)$ to the central charge of the boundary, we also compute the two- and three-point function of the holographic stress tensor $T$ and assume it to have the correct form as expected by holography. This allows us to relate the normalization of $I$ to the normalization of $T$, which in turn is related to the central charge. After computing
\be 
\langle\!\langle I\, I \rangle \! \rangle\ , \qquad 
\langle\!\langle I\, I\, I \rangle \! \rangle\ , \qquad 
\langle\!\langle T\,T \rangle \! \rangle\ , \quad\text{and}\quad 
\langle\!\langle T\,T\,T \rangle \! \rangle\ ,
\ee
one can eliminate all normalizations and compute $F(k)$ unambiguously. This then confirms eq.~\eqref{eq:F(k) c relation}.
The reader can find the details of the computation in Appendix~\ref{app:normalization}.

\subsection{A note on the torus partition function} \label{subsec:torus partition function}
Let us briefly comment on the torus partition function of global AdS$_3$. It suffers from a milder version of the same problem as the sphere, since the worldsheet path integral does not converge on a surface of genus 1. Indeed, the Ricci term in \eqref{eq:SL2R action quantum phi gamma} is missing and thus the path integral is unsuppressed in the region $\phi \to -\infty$. Instead, we can compute the one-point function on the torus of the vertex operator $I$ defined in eq.~\eqref{eq:definition I} to compute the $\mu$-derivative. Integrating with respect to $\mu$, then predicts that the string theory torus partition function takes the form
\be 
Z_{\mathbb{T}^2}=G(k) \log \mu + C_3 \sim 2 G(k) \log R+C_3\ ,
\ee
where we again reinterpret $\mu$ in terms of the size of the boundary surface as in \eqref{eq:mu R relation}. In particular, the torus partition function can compute a possible one-loop correction to the central charge.
Contrary to the sphere partition function, this now involves also the explicit form of the torus partition function of the internal CFT $X$ and thus the function $G(k)$ is background dependent. For example, in the background $\text{AdS}_3 \times \text{S}^3 \times \text{K3}$, it is known that the one-loop correction to the central charge is $+6$ and hence $G(k)=1$, see e.g.\ \cite{Beccaria:2014qea} for a direct supergravity computation. We are however not aware of a corresponding worldsheet computation to confirm this.

We learn in particular also that the central charge of the dual CFT is always one-loop exact since higher genus partition functions will never contain logarithmic contributions.

\section{Discussion}
Let us mention a few open questions.

\paragraph{Bosonic string.} One can repeat the same computation described in this paper for the bosonic string, but in this case the relation \eqref{eq:F(k) c relation} is not satisfied. Instead, one finds
\be 
F_\text{bos}(k)=\frac{c_\text{bos}(k-2)}{24k \mu}\ ,
\ee
$c_\text{bos}$ is again the boundary central charge in this case (as appearing in the OPE of the boundary stress tensor).
The fact that this does not match the expectation is perhaps not too worrying since there is no well-defined boundary CFT in this case because of the tachyon. We find it nevertheless puzzling and don't have a good explanation for the numerical value.

\paragraph{General backgrounds.} One may wonder whether this method of computation generalizes to other backgrounds. To a certain degree, the answer is yes -- we believe that one can compute all the non-analytic terms in the sphere partition function using this method in an arbitrary background. In a given background one has to identify the zero-momentum dilaton vertex operator and insertion of this operator leads to suitable derivatives of the sphere partition function. It would be very desirable to carry out this program and confirm that the sphere partition function indeed agrees with the on-shell action as computed in gravity.

However, sometimes the sphere partition function is expected to be analytic, but still well-defined, in which case this method can fail.
An example within string theory is the background of a stack of NS5-branes. The transverse directions can be described by the $\mathcal{N}=4$ cigar \cite{Callan:1991at, Giveon:1999px}, which is closely related to a coset of the $\mathrm{SL}(2,\mathbb{R})$ WZW model discussed in this paper.\footnote{For this, the NS5 branes have to be arranged in a particular circular configuration.} In this case it is reasonable to suspect that the sphere partition function computes (minus) the tension of the stack of NS5-branes, which is indeed proportional to $g_\text{s}^{-2}$ (as defined in flat space), but does not contain any logarithmic terms. As a consequence, the second derivative of the sphere partition function is actually zero and thus the tension is hidden in one of the integration constants. 
Such situations also arise in the computation of black hole entropies. For example, the Schwarzschild black hole can be $\alpha'$ corrected and embedded in string theory. The string theory sphere partition function should compute the free energy, which is fully analytic.
In such cases, there does not seem to be any known way to compute the sphere partition function directly in string theory without reducing it to a supergravity computation.

\section*{Acknowledgements}
We would like to thank Amr Ahmadain, Matt Heydeman, Shota Komatsu, Adam Levine, Raghu Mahajan, Juan Maldacena and Kostas Skenderis for discussions and Andrea Dei for comments on an early draft. We would like to thank one of the referees for pointing out a crucial sign error in the first version.
LE is supported by the grant DE-SC0009988 from the U.S. Department of Energy. SP acknowledges the support by the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award Number DE-SC0011632 and by the Walter Burke Institute for Theoretical Physics. 

\appendix

\section{Fixing the normalization} \label{app:normalization}
In this appendix, we fix the precise function $F(k)$ in \eqref{eq:sphere partition function integration constants} for the RNS superstring.
\subsection{Picture numbers}
In the main text we neglected picture numbers of the RNS superstring; or fermionic moduli space integrations in the language of supermoduli space \cite{Witten:2012bh}. In the superstring, the vertex operator \eqref{eq:definition I} sits in a supermultiplet
\be 
I=I^{(-1,-1)}+\theta I^{(0,-1)}+\bar{\theta} I^{(-1,0)}+\theta \bar{\theta} I^{(0,0)}\ ,
\ee
where $\theta$ is a Grassmann variable. Here the superscripts refer to the picture numbers of the vertex operator. We have
\be 
I^{(0,0)}=G_{-\frac{1}{2}} \bar{G}_{-\frac{1}{2}} I^{(-1,-1)}\ ,
\ee
where $G_{-\frac{1}{2}}$ is the $\mathcal{N}=1$ supercharge on the worldsheet.
$\text{OSP}(1|2,\mathbb{C})/\mathbb{Z}_2$ has two complex fermionic directions which means that the picture number of a correlation function has to sum up to $-2$ in a correlation function. Thus in a two-point function we can choose both vertex operators to have picture number $-1$, while we need to picture raise one of the vertex operators in a three-point function. Picture raising is equivalent to integration over a single supermodulus \cite{Witten:2012bh}. It also involves additional terms with ghosts, but they do not contribute to correlation functions on the sphere. See e.g.\ \cite[Chapter 13.2]{Blumenhagen:2013fgp} for a standard discussion.

Concretely, this means that \eqref{eq:mu derivative I insertion string theory} reads more precisely
\be 
\partial_\mu^3 Z_{\mathrm{S}^2}=C_{\text{S}^2} \langle I^{(-1)}(0)\, I^{(-1)}(1)\, I^{(0)}(\infty) \rangle\ ,
\ee
where we wrote the common left and right picture number only once.
Of course we can choose the normalization of picture raising in an arbitrary way, but we will compare the normalization to the correlation function of the stress tensor, and all these normalization factors will cancel out.

\subsection{Current algebra and vertex operators}
The analytically continued $\SL(2,\RR)_{k+2}$ WZW model describing string theory on Euclidean $\mathrm{AdS}_3$ has primary vertex operators $V_j(x,z)$. Even though they are of course not holomorphic, we suppress right-moving coordinates. There are also so-called spectrally flowed vertex operators, which however we do not need for our discussion \cite{Teschner:1997ft, Maldacena:2001km, Dei:2021xgh}.
The spin $j$ can take either values in $\frac{1}{2}+i \RR$ corresponding to the principal series representation of $\mathfrak{sl}(2,\RR)$ or in the interval $\frac{1}{2}<j<\frac{k+1}{2}$ corresponding to the discrete series representation. The latter correspond to short string states in target space \cite{Maldacena:2000hw}. All states of interest to us are of the latter type.

The primary fields $V_j(x,z)$ can be defined by translating with the operators $J_0^+$ and $J_0^-$ which correspond to the translation operators on the boundary of Euclidean $\mathrm{AdS}_3$, 
\begin{equation}\label{eq:trV}
V_j(x,z):=\mathrm{e}^{xJ_0^++\bar{x}\bar J_0^+}\, V_j(0,z)\, \mathrm{e}^{-x\bar J_0^+-xJ_0^+}\ .
\end{equation} 
They have the following OPE with currents
\begin{subequations} 
\begin{align}
   J^{+}(\zeta)V_{j}(x,z)&\sim \frac{1}{\zeta-z} \, \partial_{x}V_{j}(x,z) \ , \\
     J^{3}(\zeta)V_{j}(x,z)&\sim \frac{1}{\zeta-z} \, (x\partial_{x}+j)\, V_{j}(x,z) \ , \\
 J^{-}(\zeta)V_{j}(x,z) &\sim \frac{1}{\zeta-z}\,  (x^2\partial_{x}+2xj)\, V_{j}(x,z) \ .
 \end{align}\label{eq:OPE}%
\end{subequations}
The OPE with $\bar J^a$ is analogous.

The modes of the currents and the adjoint fermions satisfy the (anti)commutation relations
\begin{subequations}
\begin{align}
[J^3_m,J^\pm_n]&=\pm J^\pm_{m+n}\ , \!& [J^+_m,J^-_n]&=k m \delta_{m+n}-2J^3_{m+n}\ ,\!\!\! & [J^3_m,J^3_n]&=-\tfrac{km}{2} \delta_{m+n}\ , \\
[J^3_m,\psi^\pm_r]&=\pm \psi^\pm_{m+r} \ , \!& [J^\pm_m,\psi^\mp_r]&=\mp 2 \psi^3_{m+r}\ , & [J^\pm_m,\psi^3_r]&=\mp \psi^\pm_{m+r}\ , \\
\{\psi^3_r,\psi^3_s\}&=-\tfrac{k}{2}\delta_{r+s}\ , \!\!\!& \{\psi^+_r,\psi^-_s\}&=k \delta_{r+s}\ .
\end{align} \label{eq:commutation relations}%
\end{subequations}
We consider the NS sector and thus $r,\, s \in \mathbb{Z}+\frac{1}{2}$.
It is also useful to define the decoupled currents
\be 
\mathcal{J}^\pm_m:=J_m^\pm \pm \frac{2}{k} (\psi^3 \psi^\pm)_m\ , \qquad \mathcal{J}^3_m:=J^3_m+\frac{1}{k} (\psi^-\psi^+)_m\ , 
\ee
which satisfy a current algebra at level $\mathfrak{sl}(2,\RR)_{k+2}$ and commute with the fermions. Thus they are the currents of the $\SL(2,\mathbb{R})_{k+2}$ WZW model that we discussed in the main text. The stress tensor and the supercharge, defining the $\mathcal{N}=1$ Virasoro algebra on the worldsheet,  are then given by
\begin{subequations}
\begin{align} 
T&=\frac{1}{2k} \big(\mathcal{J}^+ \mathcal{J}^-+\mathcal{J}^- \mathcal{J}^+ -2 \mathcal{J}^3 \mathcal{J}^3-\psi^+ \partial \psi^-- \psi^- \partial \psi^++2 \psi^3 \partial \psi^3\big)\ , \label{eq:T worldsheet}\\ 
G&=\frac{1}{k} \big(\mathcal{J}^+ \psi^-+\mathcal{J}^- \psi^+-2\mathcal{J}^3 \psi^3-\frac{2}{k} \psi^3 \psi^+ \psi^- \big)\ , \label{eq:G worldsheet}
\end{align}
\end{subequations}
where normal ordering is implied. Of course, these also need to be combined with the internal CFT of the $\mathrm{AdS}_3$ compactification in order to define a critical string worldsheet theory.

\subsection{Correlation functions}
As discussed around eq.~\eqref{eq:definition I}, we are interested in correlators of the dilaton vertex operator, which is a descendant of the spin $j=1$ vertex operator. Thus, we will need the three point function of $V_1$, which is given by
\begin{equation}
\Big\langle \prod_{i=1}^3V_{1}(x_i,z_i) \Big\rangle\\
= \frac{f_{VVV} \mu^{-2}}{|x_1-x_2|^2|x_2-x_3|^2|x_3-x_1|^2}\ . \label{eq:VVV}
\end{equation}
The structure constant $f_{VVV}$ is known explicitly, but we do not require the precise form \cite{Teschner:1997ft}. We have only spelled out its $\mu$-dependence following from \eqref{eq:KPZ scaling} explicitly, since this will be important for us. 
The correlation function is $z_i$-independent since the conformal weight of $V_1$ is zero.

The vertex operator $I^{(-1)}$ is the unique vertex operator in the theory with vanishing $J_0^3$ eigenvalue. It takes the form
\begin{equation}
I^{(-1)}:= N_I\, \psi^{-}_{-1/2} \bar{\psi}^{-}_{-1/2}V_{1} \ , \qquad I^{(0)}:=G_{-1/2}\bar{G}_{-1/2}I^{(-1)}=N_I\, \mathcal{J}^{-}_{-1}\bar{\mathcal{J}}^{-}_{-1}V_1\ ,
\end{equation}
where the second equation is a straightforward calculation using the worldsheet supercharge \eqref{eq:G worldsheet} as well as the commutation relations \eqref{eq:commutation relations}. We included an arbitrary normalization $N_I$. This is the more axiomatic way of writing the vertex operator \eqref{eq:definition I}.

It is then simple to compute the string theory two-point function. The two-point function of $V_1$ with $V_1$ involves a $\delta(0)$ and is thus divergent. This is compensated in string theory by the division of the volume of the residual M\"obius symmetry and one obtains a finite result.\footnote{There is a finite factor remaining in this cancellation, but it is universal and thus unimportant for us \cite{Kutasov:1999xu, Maldacena:2001km}.}
We thus conclude that (see eq.~\eqref{eq:second derivative sphere} for the first equality)
\begin{equation}
F(k)\mu^{-1}=\langle\!\langle I^{(-1)}(x_1)\,  I^{(-1)}(x_2)\rangle\!\rangle = k^2 C_{\text{S}^2}N_I^2 f_{VV} \mu^{-1}\ , \label{eq:string two point function I}
\end{equation}
where $f_{VV}$ is the two-point function of $V_1$ with $\delta(0)$ stripped off, which is cancelled by the M\"obius volume. 
We also included the (super)ghosts, which cancel the $z_i$-dependence of the correlation function. We again spelled out the $\mu$-dependence explicitly. The factor $k^2$ is important, it came from removing the left- and right-moving fermions, which leads to the factor of $k$ appearing in the anticommutator $\{\psi_r^+,\psi_s^-\}$, see eq.~\eqref{eq:commutation relations}. 

Let us illustrate this computation a little more in detail for the case of the three point function $\langle\!\langle I^{(-1)}(x_1) \, I^{(-1)}(x_2) \, I^{(0)}(x_3)\rangle\!\rangle $ of $I$.  We begin by evaluating the corresponding CFT 3-point correlator for these operators. We can remove the free fermion and current algebra modes as follows:
\begin{align}
&\big\langle I^{(-1)}(x_1,z_1) \, I^{(-1)}(x_2,z_2) \, I^{(0)}(x_3,z_3)\big\rangle \nonumber\\
&\qquad=N_I^3\oint_{z_1} \frac{\mathrm{d}z}{z-z_1} \, \big \langle \big(\psi^-(z)-2x_1\psi^3(z)+x_1^2\psi^+(z)\big)\big(\bar \psi^{-}_{-1/2} V_{1}\big)(x_1,z_1)\nonumber\\
&\hspace{5.6cm}\big(\psi^{-}_{-1/2}\bar \psi^{-}_{-1/2}  V_{1}\big)(x_2,z_2)\big(\mathcal{J}^{-}_{-1}\bar{\mathcal{J}}^{-}_{-1}V_1\big)(x_3,z_3) \big \rangle\\
&\qquad=\frac{N_I^3k(x_1-x_2)^2 }{z_1-z_2}\big\langle \big(\bar \psi^{-}_{-1/2} V_{1}\big)(x_1,z_1) \big(\bar \psi^{-}_{-1/2} V_{1}\big)(x_2,z_2) \big(\mathcal{J}^{-}_{-1}\bar{\mathcal{J}}^{-}_{-1}V_1\big)(x_3,z_3)\big\rangle \\
&\qquad=\frac{N_I^3k(x_1-x_2)^2 }{z_1-z_2}\oint_{z_3} \frac{\mathrm{d}z}{z-z_3}\, \big\langle \big(\bar \psi^{-}_{-1/2} V_{1}\big)(x_1,z_1) \big(\bar \psi^{-}_{-1/2} V_{1}\big)(x_2,z_2)\nonumber\\
&\hspace{4.5cm} \big(\mathcal{J}^{-}(z)-2x_3 \mathcal{J}^3(z)+x_3^2 \mathcal{J}^+(z) \big)\big(\bar{\mathcal{J}}^{-}_{-1}V_1\big)(x_3,z_3)\big\rangle\\
&\qquad=-\frac{N_I^3k(x_1-x_2)^2 }{z_1-z_2}\sum_{i=1,2}\frac{1}{z_i-z_3}\left[2(x_i-x_3)+(x_i-x_3)^2\partial_{x_i}\right] \nonumber\\
&\hspace{4cm} \big \langle  \big(\bar \psi^{-}_{-1/2} V_{1}\big)(x_1,z_1)\big(\bar \psi^{-}_{-1/2} V_{1}\big)(x_2,z_2) \big(\bar{\mathcal{J}}^{-}_{-1}V_1\big)(x_3,z_3)\big\rangle\\
&\qquad=\frac{N_I^3 k (x_1-x_2)(x_2-x_3)(x_3-x_1)}{(z_1-z_2)(z_1-z_3)}\nonumber\\
&\hspace{4cm} \big \langle  \big(\bar \psi^{-}_{-1/2} V_{1}\big)(x_1,z_1)\big(\bar \psi^{-}_{-1/2} V_{1}\big)(x_2,z_2) \big(\bar{\mathcal{J}}^{-}_{-1}V_1\big)(x_3,z_3)\big\rangle\ ,
\end{align}
where we used the $x$-dependence of the three-point function \eqref{eq:VVV} in the last step.
Now we repeat the same process to strip off the antiholomorphic modes and obtain
\begin{align}
&\langle I^{(-1)}(x_1,z_1) \, I^{(-1)}(x_2,z_2) \, I^{(0)}(x_3,z_3)\rangle\nonumber\\
&\qquad=\frac{N_I^3 k^2 |x_1-x_2|^2|x_2-x_3|^2|x_3-x_1|^2}{|z_1-z_2|^2|z_1-z_3|^2} \Big\langle \prod_{i=1}^3V_{1}(x_i,z_i) \Big\rangle \\
&\qquad=\frac{N_I^3k^2f_{VVV}\mu^{-2}}{|z_1-z_2|^2|z_2-z_3|^2}\ .
\end{align}
The superconformal ghosts cancel the $z$-dependence of this correlator and we obtain for the string theory correlator
\begin{equation}
\langle\!\langle I^{(-1)}(x_1) \, I^{(-1)}(x_2) \, I^{(0)}(x_3)\rangle\!\rangle 
= k^2N_I^3 C_{\text{S}^2}f_{VVV} \mu^{-2}\ .
\end{equation}
We have, see \eqref{eq:second derivative sphere} and \eqref{eq:third derivative sphere} 
\begin{equation}
\partial_\mu  \langle\!\langle I^{(-1)}(x_1) \, I^{(-1)}(x_2)\rangle \! \rangle =\langle\!\langle I^{(-1)}(x_1) \, I^{(-1)}(x_2)\,I^{(0)}(x_3) \rangle \! \rangle\,,
\end{equation}
which gives
 \begin{equation}
  N_If_{VVV}=-f_{VV}\ . \label{eq:c3SS}
 \end{equation}

This of course does not completely determine the function $F(k)$ appearing in \eqref{eq:second derivative sphere} and \eqref{eq:third derivative sphere}. We will fix it by computing the correlation function of the stress energy tensor on the boundary.
The operator corresponding to the stress tensor on the boundary is the unique physical operator with $J_0^3$ eigenvalue 2 and $\bar{J}_0^3$ eigenvalue 0 (and is thus a holomorphic field of conformal weight $(2,0)$ on the boundary). It takes the form
 \begin{align}
T^{(-1)}&:=N_T\, \big(\psi^+_{-1/2} - J_0^{+}\psi^{3}_{-1/2}+\frac{1}{2}(J_{0}^+)^2\psi^{-}_{-1/2}\big)\, \bar \psi^{-}_{-1/2}V_{1}\\
&=3N_T\, \big(\psi^+_{-1/2} - \psi^{3}_{-1/2}J_0^{+}+\frac{1}{2}\psi^{-}_{-1/2}(J_{0}^+)^2\big)\, \bar \psi^{-}_{-1/2}V_{1}\ .
\end{align}
The corresponding operator in picture 0 takes the form
\begin{equation}
T^{(0)}:=G_{-1/2}T^{(-1)}=N_T\, \big(J^+_{-1} - J_0^{+}J_{-1}^3+\frac{1}{2}(J_{0}^+)^2J_{-1}^{-}\big) \, \bar{\mathcal{J}}^{-}_{-1}V_{1}\ .
\end{equation}
We compute
\begin{equation}
 \langle\!\langle T^{(-1)}(x_1) \,  T^{(-1)}(x_2)\rangle \! \rangle =\frac{3k^2N_T^2C_{\text{S}^2} f_{VV}\mu^{-1}}{(x_1-x_2)^4}\ ,\label{eq:2pointTSS}
\end{equation}
where we again allowed for an arbitrary normalization for the vertex operator $T$.
From the perspective of the dual CFT with central charge $c$,  we have 
\be\label{eq:2pointTdualSS}
 \langle\!\langle T^{(-1)}(x_1) \,  T^{(-1)}(x_2)  \rangle\! \rangle =\frac{c}{2(x_1-x_2)^4}\ ,
 \ee
 Comparison of eqs.~\eqref{eq:2pointTSS} and \eqref{eq:2pointTdualSS} tyields
\begin{equation}\label{eq:c1SS}
6k^2N_T^2 C_{\text{S}^2} f_{VV}\mu^{-1}=c\ .
\end{equation}
We finally compute
 \begin{equation}
 \langle\!\langle T^{(-1)} (x_1)\,  T^{(-1)}(x_2) \, T^{(0)}(x_3)\rangle \! \rangle 
= \frac{6k^2N_T^3 C_{\text{S}^2}f_{VVV}\mu^{-2}}{(x_1-x_2)^2(x_2-x_3)^2(x_3-x_1)^2}\ .\label{eq:3pointTSS}
\end{equation}
Comparing with the generic three-point function of the stress tensor in a CFT,
  \begin{equation}
 \langle\!\langle T^{(-1)} (x_1)\,  T^{(-1)}(x_2) \, T^{(0)}(x_3)\rangle \! \rangle 
=\frac{c}{(x_1-x_2)^2(x_2-x_3)^2(x_3-x_1)^2}\ ,\label{eq:3pointTdualSS}
\end{equation}
we obtain the equation
\be\label{eq:c2SS}
6k^2 N_T^3 C_{\text{S}^2} f_{VVV}\mu^{-2}=c\ .
\ee
We have now obtained the three equations \eqref{eq:c3SS}, \eqref{eq:c1SS} and \eqref{eq:c2SS},
which allow us to compute the function $F(k)$ appearing in the string sphere partition function \eqref{eq:sphere partition function integration constants}. From \eqref{eq:string two point function I}, we have
\begin{align}
F(k)=k^2 N_I^2C_{\text{S}^2} f_{VV} =\frac{k^2 C_{\text{S}^2}f_{VV}^3}{f_{VVV}^2}=\frac{c}{6\mu}\ , \label{eq:F(k) in terms of c}
\end{align}
where we used \eqref{eq:c3SS} in the first equality and the corresponding ratio of \eqref{eq:c1SS} and \eqref{eq:c2SS} in the second. This is the expected result as explained after eq.~\!\eqref{eq:F(k) c relation}.

\bibliographystyle{JHEP}
\bibliography{bib}
\end{document}